OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = mc012-1.jpgx + 7; (3, 11)

OpenStudy (anonymous):

y=2x+7 ; (3,110

OpenStudy (anonymous):

If it passes through (3,11), and its parallel to that line, then it has to have the SAME slope as the original equation, but a different intercept. We can find that intercept with the point you give us. Using the equation form: Y-y1= m(x-x1) we can solve for Y-11=2(x-3) (Some teachers will take this as an answer, but others might want point slope) Y=2x -6 + 11 Y=2x + 5